OpenStudy (anonymous):

During a game of hearts, three cards are dealt from a standard deck of cards, one at a time without replacement, find the probability thats: A) all are kings? B) all are red cards?

hartnn (hartnn):

how many total cards are there and out of that how many kings are there?

OpenStudy (anonymous):

4/52 reduced to 1/13

hartnn (hartnn):

that is the probability of getting one king from 1 deck....

hartnn (hartnn):

but here u are dealing 3 cards, so for the first card to be king the probability = 4/52=1/13 as u said now there are 51 cards and 3 kings,so the probability that 2nd card dealt is also king is 3/51 similarly,the 3rd card dealt is also king is 2/50 so overall probability is 4/52*3/51*2/50 got this first part??

OpenStudy (anonymous):

ok i got it makes sense

hartnn (hartnn):

still want 2nd part?

OpenStudy (anonymous):

yeah i am not good with probability

OpenStudy (anonymous):

13/52 for reds?

hartnn (hartnn):

that is the probability of choosing 1 red from 52 cards but here u are dealing 3 cards, so for the first card to be red the probability = 13/52=1/4 as u said now there are 51 cards and 12 red cards remaining,so the probability that 2nd card dealt is also red is 12/51 similarly,the 3rd card dealt is also red is 11/50 so overall probability is (13/52)*(12/51)*(11/50) got this second part??

OpenStudy (anonymous):

i get it now i appreciate the help my friend.